The expanded form of the number 9578 is 
 The expanded form of the number 9578 is 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Consider the given number 9578,
The place value of 8 is ones = 8 × 1
The place value of 7 is tens = 7 × 10
The place value of 5 is thousand = 5 × 100
The place value of 9 is ten thousand = 9 × 1000

9 × 1000 + 5 × 100 + 7 × 10 + 8 × 1
is the answer

Thanks